OpenStudy (anonymous):

The molar mass of a acid is 40 . 50 ml of one normal Ca(OH)2 is neutralized by 0.5g of acid The basicity of the acid is ?

OpenStudy (anonymous):

@ScottWu @JFraser

OpenStudy (anonymous):

"The basicity of the acid is ?" sorry i'm a little confused

OpenStudy (anonymous):

Basicity = No of H+ ions

OpenStudy (anonymous):

sorry, what is "one normal..."?

OpenStudy (anonymous):

n * (0.5/40) = 2 * (weight of Ca(OH)2 / 74)

OpenStudy (anonymous):

that's what i've got...

OpenStudy (anonymous):

or this way n* (0.5/40) = 1(N)/2 *0.005

OpenStudy (anonymous):

Me too confused no prob just look my next prob

OpenStudy (anonymous):

Since I wonder what is "one normal"... (my mother tongue is not English, I might learn about it while I missed in the gap of language.)